KyledaGreat:

The half-life of silver-105 is approximately 41.3 days. Step 2 of 3 : How much of a 2 gram sample of silver-105 would remain after 86 days? Round to three decimal places. ___________________ grams

Vocaloid:

using the equation from part 1 A(t) = A0 * a^t, where we calculated a = 0.983357 A0 is 2 grams t is 86 days so plugging things in: A(t) = 2 * 0.983357^86

KyledaGreat:

0.47227436

Vocaloid:

good, but since it says to round to 3 decimal places, just 0.472

KyledaGreat:

okay